OpenStudy (anonymous):

Determine algebraically whether the graph is symmetric with respect to the x-axis, the y-axis, and the origin. 5) x2 + 2y4 = 5

OpenStudy (amistre64):

if we solve for y .....

OpenStudy (amistre64):

the function looks to me to be even overall so wrtO seems out of the picture

OpenStudy (amistre64):

x2 + 2y4 = 5 2y4 = 5-x^2 y = +- 4rt((5-x^2)/2)

OpenStudy (amistre64):

if x=1 is the same as x=-1; theres a high chance its sym about the y axis

OpenStudy (anonymous):

as im getting "y" by itself how do I get rid of the power of 4?

OpenStudy (amistre64):

take the 4throot

OpenStudy (amistre64):

to undo a square, take the sqrt to undo a cube, take the cbrt to undo the nth, take the nroot th

OpenStudy (amistre64):

hmmm, now that i think of it more, lets take this example